Kyky232:

Find the perimeter of the following shape, rounded to the nearest tenth: (pic in replies) 12.5 15.7 16 16.5

Kyky232:

1 attachment
AZ:

Do you know the distance formula?

Kyky232:

no...

AZ:

Have you seen this formula before? \( \text{distance} = \sqrt{(x_2 - x_1)^2 + (y_2-y_1)^2} \)

Kyky232:

nope/

AZ:

mhmm have you learned about special triangles then perhaps? Such as the 45-45-90 degree triangle Or Pythagoras theorem? a^2 + b^2 = c^2

Kyky232:

I know all sides angles equal 180 0-0

Kyky232:

thas bout it

AZ:

To find the perimeter, we have to find the lengths of each side and then add them up

AZ:

What are the coordinates of point A?

Kyky232:

(5,5)

AZ:

Good! What about point B?

Kyky232:

(7,3)?

AZ:

Perfect! Now to find the length of that side that's between point A and B, we use the distance formula \( d = \sqrt{(x_2 - x_1)^2 + (y_2-y_1)^2}\) where you have two points \( (x_1, y_1)\) and \( (x_2, y_2)\)

yaboiaction:

@AZ

AZ:

So we have the two points you mentioned (5, 5) and (7, 3) it's in the form of \( (x_1, y_1)\) and \( (x_2, y_2)\) So let's plug in the numbers and simplify, okay?

Kyky232:

it would be 2.83?

Kyky232:

well if you simplify

AZ:

\( d = \sqrt{(7- 5)^2 + (3-5)^2} \)

AZ:

Good! That's one of the sides of our rectangle.

AZ:

What are the points C and D?

Kyky232:

another side would be 5

AZ:

Yes!!! You can also see that if you drew a triangle with that side as the hypotenuse, that it would form a 3, 4, 5 triangle

Kyky232:

so then you add those and multiply by 2 because its a rectangle right?

AZ:

So now we have 

AZ:

Yes! Exactly!

Kyky232:

15.66??

Kyky232:

theres only 15.7 on there. So I think im supposed to round

AZ:

That's your answer!

Kyky232:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ That's your answer! \(\color{#0cbb34}{\text{End of Quote}}\) thank you!!!

AZ:

It was my pleasure! :)